OpenStudy (anonymous):

Solve 3x^2 - 6x = -3

OpenStudy (anonymous):

X=1 X=1/3

OpenStudy (anonymous):

actually forget the x=1/3

OpenStudy (anonymous):

3x^2 - 6x = -3 Add 3 to both sides 3x^2 - 6x + 3 = 0 Factor out 3 3(X^2 - 2x +1) = 0 Divide both sides by 3 x^2 - 2x + 1 (x-1)(x-1) x-1 = 0 x=1